In a Math test, the mean score is 45 and the standard deviation is 4. Assuming normality, what is the probability that a score picked at random will lie below 38


1
Expert's answer
2021-03-29T09:56:13-0400

X ~ N(45, 42)

"P(X<38) = P(\frac{x-μ}{σ} < \frac{38-45}{4}) \\

= P(Z < -1.75) \\

= 0.040 \\

= 4.0 \%"